OpenStudy (anonymous):

If x is a positive angle smaller than 90° and cos(x) = 1/3, then sin(x) =

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

sinx = sqrt(1 - cos^2x) = sqrt(8/9) = 2sqrt2 / 3

OpenStudy (anonymous):

ok - i used the trig identity sin^2x + cos^x = 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no probs - plz click on good answer to close the question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ty

OpenStudy (id21):

jimmyrep great! :)

OpenStudy (anonymous):

yes he is

myininaya (myininaya):

cosx=1/3=adjacent/hypenuese so the opposite side can be found be Pythagorean thm 1^2+(opposite)^2=3^2 1+(opposite)^2=9 (opposite)^2=9-1 (opposite)^2=8 opposite=sqrt{8} sinx=opposite/hypenuese=sqrt{8}/3